# A cross-platform analysis of polarization and echo chambers in climate change discussions

### v2

@@ Abstract
With the intensification of climate change discussion, social media has become prominent in disseminating reliable and unreliable content. In this study, we present a cross-platform analysis on Youtube and Twitter, and examine the polarization and echo chambers in social media discussions in four datasets related to climate change: COP27, IPCC, Climate Refugees, and Doñana. We have identified communities of users spreading misinformation on Twitter, although they remain relatively isolated from the rest of the network. The analysis by interaction type reveals that climate change sceptics use mentions to draw the attention of other communities. The YouTube posts referenced on Twitter reveal a strong correlation in the community organisation of social media, suggesting a platform alignment. Moreover, we report the presence of echo chambers in YouTube post-sharing related to mainstream and sceptical content.
While a certain level of societal conflict can benefit democracies, an excessively polarized social environment can detrimentally affect their self-governing capacity [1], especially when it pertains to issues that require a high level of consensus to be addressed. The convolution of a highly polarized social environment and an overload of fast-spreading controversial or unreliable information impacts public perception of reality [2]. This dynamic, exacerbated during emergencies and crises [3, 4], further complicates the quest for solutions, rendering resolution increasingly challenging.
A social system is considered to be maximally polarized when the population is divided into two groups of similar size with diametrically opposed views that remain almost isolated [5, 6]. Studies and experiments across various platforms [7, 8, 9] have shown an increase in polarization attributed to the use of social media. However, the precise impact of online social networks on public debate remains ambiguous [10]. Evidence might vary across topics of discussion and online platforms. While Twitter tends to display high polarization, other platforms, such as WhatsApp, have even shown de-polarization [11]. The analysis of Twitter debates across different topics [12] revealed that the discussion on Obamacare was far less polarized than the debate on abortion. Other platforms such as YouTube or Instagram also display polarization [13, 14, 15, 16] Additionally, external events have been observed to heighten activity and polarization on social media platforms [17, 18, 19].
Even though the interest in measuring social and political polarization dates from before the digital era [20, 6], the widespread of online social platforms and data accessibility sparked the development of measures to quantify it. There are two main categories of polarization metrics, along with combinations of them, employed in the analysis of digital traces within online social platforms: those centred on textual analysis [21], and those focusing on interactions and network structure [22].
The first metrics developed to quantify the macroscopic level of structural polarization primarily concentrate on the modularity and the analysis of community boundaries [23, 24, 22]. However, these metrics only provide a partial perspective on polarization, as they largely disregard the multi-scalar organisation of networks and the role of individual nodes. More refined metrics, which capture long-range structural properties, include label propagation [25], centrality measures, and random walk diffusion [26]. Recent studies have aimed to integrate the analysis of structural polarization with individual opinions [12]. Even though most of the measures assume a dipole organisation, there have been recent developments to quantify multipole polarization [27].
Several studies indicate that a phenomenon closely intertwined with the rise of polarization in social media is the prevalence of echo chambers [28, 29, 30], where interactions predominantly occur among individuals who share similar beliefs and ideas. Research exploring the existence and strength of echo chambers across various social media platforms has revealed a spectrum of prevalence and intensity [29, 7, 31, 14, 32, 33]. Notably, platforms like Facebook and Twitter exhibit more pronounced echo chambers compared to Gab and Reddit. The presence of echo chambers can be quantified through the political leaning of an individual and its neighbourhood [34]. As interest in modelling the dynamics of polarization and echo chambers grows [35, 36], efforts to discern the multiple factors influencing social media interaction patterns face significant challenges. Even the recommendation algorithms employed by these platforms can impact observed dynamics [37, 38, 39]. For instance, platforms may inadvertently exacerbate polarization by limiting exposure to opposing viewpoints in users’ feeds [8].
@@
The combination of a polarized environment [40] and psychological traits [41], such as confirmation bias [42, 43, 44], has been closely associated with the dissemination of misinformation [45]. This spread of unreliable and fake news significantly impacts society, influencing outcomes such as presidential elections [46, 47] and vaccine hesitancy [48, 49]. Despite these challenges, there have been concerted efforts to combat the spread of fake news [50]. Nonetheless, debates persist regarding the distinction between reliable and unreliable news cascades. Although initial studies highlighted structural differences [51], subsequent analyses have suggested that disparities may be more attributable to differences in cascade size [52].
Besides politics, climate change stands out among the topics in which unreliable news has proliferated [53, 54]. The existence of climate change has been proven repeatedly from a scientific perspective [55, 56, 57, 58] as has its anthropogenic contribution [59, 60, 61]. There are multiple pieces of evidence on the impact of climate change on health [62], the economy [63], the agriculture [64] or the overall ecological system [65], among others. However, there has been an increasing polarization in social media around the impact of climate change and its anthropogenic nature [66]. Previous research has focused on the origins of this polarization [67], its relation to political affiliations [68], and the influential roles played by media outlets [69] and corporations [70]. Recently, an increasing polarization was reported in locations suffering from extreme weather events [71].
In this work, we analyse the climate change discussion in four datasets from Twitter and YouTube across different contexts. The degree of polarization varies across datasets due to their distinct characteristics and is higher in the Doñana dataset, possibly due to its close relation with local politics. We analysed the interaction between the two largest communities spreading reliable and unreliable content and confirmed a weak interaction between them. Additionally, we provide a cross-platform perspective by examining YouTube links shared on Twitter (currently X) and analysing the bipartite graph derived from these links. Our findings reveal a connection between communities across both platforms, indicating a strong interconnection across social media platforms.
@@ I Data Description
We extracted four datasets related to climate change topics from Twitter through the Twitter Academic API. The first one is related to the 2022 United Nations Climate Change Conference (COP27), the second to the release of the 6th Assessment Report by the Intergovernmental Panel on Climate Change (IPCC), the third to the climate refugee crisis resulting from extreme weather conditions (Climate Refugees), and the fourth to a new legislation concerning the natural reserve of Doñana in the south of Spain (Doñana). The tweet recollection was conducted with Twarc [72] by searching for keywords related to each discussion (Appendix A).
Network Minimum date Maximum date Number of users Tweets Retweets Replies Quotes
COP27 2022-09-01 2022-11-27 1351903 866753 4977874 205973 175078
IPCC 2023-03-18 2023-03-26 157056 31138 267971 45863 7751
Clim. Refugees 2008-03-10 2022-12-31 841454 384267 1376057 139699 38909
Doñana 2019-01-01 2023-04-30 290782 139478 1187646 135245 25056
@@
In Table 1 we present the main statistics of the datasets and in Fig. S1 the daily tweet activity, where the Climate Refugees and Doñana datasets exhibit longer timespans than the IPCC and COP27 datasets. The IPCC dataset covers only nine days, coinciding with the heightened discussion surrounding the release of the 6th Assessment Report. We provide in Appendix B the most central users based on the total inflow of retweets (Tables S1, S2, S3, and S4) and the most frequent languages in the ten largest communities of each dataset (Tables S5, S6, S7, and S8).
We extracted the corresponding YouTube dataset by analysing the URLs referenced in each Twitter dataset and downloading the posts and comments using the YouTube Data API. To ensure relevance, we filtered the posts by searching for occurrences of certain keywords in each video title or description (Appendix A).
Network Videos Min. Date Max. Date Avg. views Avg. likes Avg. comments
COP27 624 2008-02-11 2023-05-12 571964.35 5860.32 306.99
IPCC 145 2007-11-03 2023-06-09 350178.66 8991.67 1455.16
Clim. Refugees 215 2009-11-22 2023-11-06 243720.27 3604.81 527.40
Doñana 191 2009-01-11 2023-06-01 101098.32 3622.68 246.25
@@
We present a summary of the YouTube datasets in Table 2, which includes the number of videos, the minimum and maximum dates, and their average statistics. Unlike the Twitter dataset, the period of the comments is longer due to users frequently referencing older videos. Fig. S2 illustrates the time series of comments in each dataset. Each dataset exhibits peaks at different times based on its unique characteristics. In particular, we observe a strong alignment in the IPCC and COP27 datasets with the event dates.
II Twitter analysis
@@
We evaluated the global polarization on Twitter by considering the unweighted network of retweets where links are undirected and indicate a retweet interaction between two users. We used the Metis partition algorithm [73] to split the graph into two clusters, and we calculated the modularity [74, 75], the E-I index [76] and the adaptive E-I index [77] which accounts for the uneven size of clusters. The results for the four datasets analysed are presented in Fig. 1. Following the methodology proposed in [22], we computed polarization metrics in the observed network $\Phi(G)$, as well as in the ensemble of configuration models with equivalent degree distribution $\Phi(G_{CM})$ [78], and calculated the denoised value $\hat{\Phi}(G)=\Phi(G)-\Phi(G_{CM})$.
@@
Figure 1: Polarization in the Twitter dataset. Quantification of the polarization in Twitter: (a) modularity, (b) E-I index, and (c) adaptive E-I index. We provide the measures for the observed networks in red, across 100 realizations of the configuration model with preserved degree sequences in blue, and the difference between them (denoised value) in yellow.
@@
All networks exhibit polarization according to the denoised value. In particular, the COP27 and Doñana datasets displayed larger denoised values, suggesting a stronger polarization around the topics. The higher polarization observed in the COP27 dataset may be attributed to the presence of communities using different languages that do not interact with each other (Table S5). The Doñana dataset captures the discussion around a new law related to natural park preservation, which involves political parties and users with strong political affiliations. The lower polarization observed in the IPCC dataset may be due to its nature as a scientific report release rather than an international conference, potentially attracting a more scientifically educated audience and inducing a lower level of polarization. The standardised values [22] (Fig. S3) indicate that polarization is stronger in the Climate Refugees dataset.
We examine the network of retweets and mentions between the $10$ largest communities of the networks in Fig. 2. We identified these communities using modularity optimisation on the directed weighted graph of retweets [75]. The interaction patterns between the largest community (labelled as 0) and the remaining communities vary across datasets and types of interactions. In the IPCC and COP27 networks, the smaller communities mostly retweet content from the largest community, which is more isolated and engages in fewer interactions with the rest of the network. Those largest communities feature official accounts such as the Secretary General of the United Nations or the Intergovernmental Panel on Climate Change (Tables S1 and S2). Conversely, the Doñana network exhibits greater flow heterogeneity between communities, with community 5 playing a central role in attracting most of the flow. The detailed overview of the users in the community reveals the presence of the president of Spain and the ruling political party (Table S4). There are significant structural differences depending on the type of interaction. For instance, in the IPCC network, there are relatively few retweets from community 1 to community 0, but there is a high volume of mentions.
@@
Figure 2: Community interaction network. Interaction between the ten largest communities in (a) COP27, (b) IPCC, (c) Climate Refugees, (d) Doñana. Arrows indicate the direction of the interaction, the colour and width, and the volume. The size of the dots represents community size. In red, we show the interaction through retweets and, in blue, through mentions.
To assess the political bias and reliability of each community, we calculated the user values by averaging over the URLS they shared. Information on the political bias and reliability of news outlets was obtained from MediaBias fact-check [79], where positive and negative political bias correspond to right-wing and left-wing media, respectively. Outlets with higher reliability are considered more trustworthy and tend to share more verified news. For the Doñana dataset, the reliability and bias were based on a separate dataset [80], given that most media sources are Spanish. To facilitate the analysis, we have normalised the bias within the range [-1, 1] and the reliability within the range [0, 1]. The minimum and maximum bias are set according to the largest absolute value observed. In Fig. 3(a)-(d), we show the density plot of community reliability as a function of their political bias according to the shared URLs in the IPCC, COP27, Climate Refugees, and Doñana networks. We computed user-level values by averaging over the sources they shared, and then calculated community-level values by averaging over the users. All datasets have a left bias as most communities display negative values, particularly pronounced in the IPCC and Climate Refugee networks. Possible reasons for this include bias in the source data, climate change sceptics sharing non-official sources, or an over-representation of left-biased users in climate change discussions. Left-wing users might interact more in the context of climate change due to their higher awareness. The left bias is particularly pronounced in the IPCC and Climate Refugee networks, where the cluster of large communities is centred around negative values. However, the bias is more dispersed in the Doñana dataset, possibly due to differences in the shared media sources or a wider ideological variety. We show the user distributions of bias and reliability for the seven largest communities in Fig. 3(e)-(h) with most of them peaked at negative values (left bias). The main communities with right bias we have identified are community one in the COP27 and IPCC, community six in Climate Refugees, and community zero in Doñana.
@@
Figure 3: Reliability as a function of the political bias in the networks studied. (a-d) Density plot of the distribution of reliability as a function of the bias for the communities detected in (a) COP27, (b) IPCC, (c) Climate Refugees and (d) Doñana. (e-h) Distribution of bias and reliability in the seven largest communities for (e) COP27, (f) IPCC, (g) Climate Refugees and (h) Doñana.
Based on the previous findings and manual inspection of the communities (See Tables S1, S2, S3, and S4), we identified the two largest communities with opposing political leanings. In the IPCC and COP27 datasets, community 0 represents the largest community aligned with the mainstream climate change narrative, while community 1 is the primary advocate of a counter-narrative message. However, in the Doñana network, the labels are swapped because the right-biased community has a larger user base. For the Climate Refugees dataset, most of the top 10 communities exhibit similar bias-reliability values, except for community 6, which stands out with low reliability and a right bias. In this dataset, we designate community 1 as the primary left-biased community, as it includes well-known influencers such as Greta Thunberg and engages in more interactions with the sceptic community.

### v3

# Multifaceted polarisation and information reliability in climate change discussions on social media platforms

@@ Abstract
Social media platforms like YouTube and Twitter play a key role in disseminating both reliable and unreliable information about climate change. This study analyses the topology of interactions in Twitter and their relation to cross-platform sharing, content discussions and emotional responses. We examined climate change discussions across four topics: the 27th United Nations Climate Change Conference, the Sixth Assessment Report of the United Nations Intergovernmental Panel on Climate Change, climate refugees, and Doñana Natural Park. While retweets reinforce in-group cohesion in the form of echo chambers, inter-group exposure is significant through mentions, suggesting that exposure to opposing views intensifies polarisation, rather than mitigates it. Ideological divides feature content differences accompanied by steeper negative sentiments, especially from right-leaning communities prone to share low-reliability information. We identified a topological alignment between platforms, indicating that ideological communities span multiple sites. Our findings show that climate change polarisation is multifaceted, involving ideological divides, structural isolation, and emotional engagement. These results suggest that effective climate policy discussions must address the emotional and identity-driven nature of public discourse and seek strategies to bridge ideological divides.
Polarisation has become a defining feature of contemporary political and social life. While a certain level of conflict and disagreement is essential to the functioning of democracies, an excessively polarised social environment can detrimentally affect their self-governing capacity [1], particularly when it pertains to issues, like climate change, that require broad consensus and coordinated action. In this paper, we examine polarisation in climate change discourse on Twitter (currently X) and YouTube, using a multidimensional approach. To capture different social and geographical contexts, we analyse four datasets related to milestone events: the 2022 UN Climate Change Conference (COP27), the release of the 6th Assessment Report by the Intergovernmental Panel on Climate Change (IPCC), the emergent issue of climate refugees, and a politically charged local debate in Spain over environmental regulation in Doñana.
The term polarisation is used across disciplines to describe the division of a population into two well-separated and opposing groups. It can be understood both as a state – the extent to which opinions or characteristics are opposed at a given time – and as a process – the movement toward greater extremes over time [2]. Furthermore, polarisation has a specific meaning related to the distribution of opinions, beliefs, attitudes, or other measurable attributes in a population (such as wealth or income) [3]. Operationalisations of this concept emphasise the formation of distinct groups in the distribution, with high homogeneity within the group and heterogeneity between groups. The core notion being that polarisation is high when a distribution is multi-modal and maximal when it is bimodal (i.e., when there are multiple distinct ’peaks’ corresponding to social groups) rather than a continuum [4].
Esteban and May [3] frames polarisation in terms of ’identification’ and ’alienation’. Polarisation is conceptualised as the result of people strongly identifying with others in their own group while feeling alienation (distance or hostility) toward those in other groups. A formal model is proposed, where society is more polarised when it contains relatively cohesive subgroups that are far apart from each other. High identification within groups means that each of the groups of individuals has internal similarity or solidarity, and high alienation means there is strong effective distance or antagonism between those groups. This idea leads to well-defined mathematical measures in which polarisation is essentially the weighted sum of pairwise ’antagonisms’ between individuals, increasing when people cluster into separated groups.
The shift of political views towards extreme positions and the formation of separate cohesive groups often is related to the presence of echo chambers, where interactions predominantly occur among individuals who share similar beliefs and ideas [5, 6, 7]. As interest in measuring [6, 8, 9, 10, 11] and modelling [12] the dynamics of polarisation and echo chambers grows, efforts to identify the factors influencing the interaction patterns face significant challenges. The recommendation algorithms employed by these platforms influence these dynamics [13, 14, 15] by limiting exposure to opposing viewpoints in users’ feeds [16]. Yet the rise of polarisation is not only an effect of homophily as it has been shown that individuals are often exposed to opposed opinions [17].
While polarisation is a natural aspect in pluralistic societies, its intensification over the last decades is a matter of debate both among academics and in public discourse [18, 2, 19, 20]. Multiple studies point to the fact that the widespread adoption of social media has driven an increase in polarisation as observed across experimental settings [21] and various platforms [8, 16, 22]. However, its presence and magnitude vary across subjects, can be triggered by external events [23, 24, 25, 26] and is steeper around political discussions than in more general subjects [27]. Platforms also play a role, while polarisation has been observed on Twitter, YouTube, or Instagram [28, 29, 30, 31], in others, such as WhatsApp, there is a depolarisation effect [32]. Even though the interest in measuring social and political polarisation dates from before the digital era [33], the widespread use of online social media and data accessibility sparked the development of frameworks to quantify it. The techniques developed vary depending on the perspective, while the topology of interactions provides insights into the preferences to form social bonds based on political views [34], textual analysis can capture the divergence in emotions and policies [35]
@@
The combination of a polarised environment [36] and behavioural factors, such as confirmation bias [37, 38], has been closely associated with the dissemination of low-reliability information [39], influencing the outcome of presidential elections [40] and increasing vaccine hesitancy [41]. Despite these challenges, there have been concerted efforts to combat the spread of fake news [42]. Nonetheless, debates persist regarding the distinction between reliable and unreliable news cascades. Although initial studies highlighted structural differences [43], subsequent analyses have suggested that disparities may be more attributable to differences in cascade size [44].
Polarisation manifests across political issues, but also other kinds of events and subjects as we have witnessed an increase in climate change discussions and related events [45, 46, 20, 26]. Previous research has focused on the origins of this polarisation [47], its relation to political affiliations [48], and the influential roles played by media outlets [49] and corporations [50]. Recently, an increasing polarisation has been reported in locations affected by extreme weather events or environmental disputes [46]. The sentiment towards climate change mitigation measures varies significantly across regions, shaped by local political and cultural contexts [51]. While the spread of low-reliability information in these contexts has been noted, its connection with polarised environments remains unclear [52, 53, 54]. Our study addresses this gap by studying a highly localised and politically charged debate over environmental protection in Spain, Doñana, where the interplay between misinformation, partisanity, and institutional trust is especially salient.
The broad concept of polarisation can then be further specified, with particular relevance to our work, by focusing on the notions of political and affective polarisation. Political polarisation usually refers to the distance between political preferences, identities, or parties within a society [2]. In relation to this, the analysis of a retweet network relies on the empirical and theoretically supported assumption that the way users share information is related to and a source of their collective political identity [55, 56]. In the same line, information sharing is related to the practice of networked gatekeeping that put in relation elite and mass polarisation [57] and lead to the formation of affective publics [58]. Affective polarisation is also emphasised in political science and is distinct from policy disagreement. It describes the phenomenon of partisans feeling not just different, but actively antagonistic toward each other. [19, 59]. In other words, it’s the polarisation of feelings or attitudes toward the social groups representing the other, rather than toward policy issues themselves. This is well captured through expressions of antagonism between groups, such as emotional tone in communication. In relation with the structural polarisation metrics we apply in this work, it is central to the notion of how users use retweet and mention as different ways of signalling in- and out-group relations and identities [60].
The initial metrics to quantify the macroscopic level of structural polarisation primarily concentrate on modularity and the analysis of community boundaries [61, 62, 34]. However, those only provide a partial perspective on the organisation of social connections, as they largely disregard the multi-scalar organisation of networks and the role of individual nodes. Metrics designed to capture long-range structural properties include label propagation [63], centrality measures, and random walk diffusion [64]. Even though most assume a dipole organisation, recent developments can capture multipole polarisation [65]. Structural measures in dealing with the networked organisation of social relations [66] provide a measure of polarisation being content-agnostic. Natural language processing techniques are essential for assessing the content of the divergence on policy-related issues [32, 45] and evaluating the emotions of individuals and their interactions [60]. While our approach focuses on structural and affective dynamics at a group level, recent studies have aimed to integrate these with individual-level opinions [31] providing further context into the structural organisation of different ideological perspectives.
Building on previous literature that tends to isolate structural, ideological, or affective polarisation, this study offers an integrated approach to capture their intersection in climate change discourse. Drawing on Twitter and YouTube data, we examine how community structure, emotional tone, and media reliability interact across platforms and contexts. To do so, we combine network analysis (to assess structural patterns), ideological measures (using political bias and source reliability), and sentiment metrics (focused on toxicity in quote tweets and cross-group mentions). This multidimensional approach allows us to observe how different dimensions of polarisation reinforce one another. Users with similar ideological leanings tend to cluster in retweet networks and express aligned affective reactions, particularly in politically salient debates. Interactions across communities, while rare in the form of retweets, are more frequent through mentions and often carry antagonistic sentiment—suggesting emotional confrontation despite low structural connectivity.
Our main contribution is the development of an integrative framework that captures structural, ideological, and affective polarisation across digital platforms. Applying this framework to four climate-related events, we examine how polarisation manifests differently depending on the topic and geographical context, ranging from global summits to locally contentious environmental debates. This multidimensional analysis reveals consistent links between the structure of online communities, the emotional tone of user interactions, and the reliability of the content being shared. Notably, we observe that polarisation is not confined to a single platform: ideologically aligned groups on Twitter also engage with similar content on YouTube, indicating a broader cross-platform coherence. These findings underscore the importance of moving beyond single-platform or single-dimension studies in order to understand the full complexity of polarisation in digital climate discourse.
@@ I Data Description
We extracted four datasets related to climate change topics from Twitter through the Twitter Academic API. These datasets correspond to the 2022 United Nations Climate Change Conference (COP27), the release of the 6th Assessment Report by the Intergovernmental Panel on Climate Change (IPCC), the climate refugee crisis resulting from extreme weather conditions (Climate Refugees), and to legislation affecting the natural reserve of Doñana in the south of Spain (Doñana). Tweets were collected using Twarc [67] by searching for keywords related to each discussion. The keywords selected were closely related to each event. For COP27 the keywords included COP27 and related concepts such as LossAndDamage. For IPCC the keywords were IPCC and its official account (@IPCC_CH). For climate refugees the keywords were related combinations included climate refugee(s) and climate migration. For Doñana only the keyword Doñana was used. The datasets for specific events (COP27 and IPCC) were filtered within the span of the pre and post-event dates. The Climate Refugees and Doñana datasets span a longer period compared to the IPCC and COP27 datasets. The IPCC dataset, in particular, covers only nine days, coinciding with the peak discussion surrounding the release of the 6th Assessment Report (See Table 1). We extracted the corresponding YouTube dataset by analysing the URLs referenced in each Twitter dataset and downloading the posts and comments using the YouTube Data API. To ensure relevance, we filtered the posts by searching for occurrences of specific keywords in each video title or description (Section S1).
Min. date Max. date Users Tweets Retweets Replies Quotes
COP27 2022-09-01 2022-11-27 1,351,903 866,753 4,977,874 205,973 175,078
IPCC 2023-03-18 2023-03-26 157,056 31,138 267,971 45,863 7,751
C. Refugees 2008-03-10 2022-12-31 841,454 384,267 1,376,057 139,699 38,909
Doñana 2019-01-01 2023-04-30 290,782 139,478 1,187,646 135,245 25,056
@@ C. Refugees 2008-03-10 2022-12-31 841,454 384,267 1,376,057 139,699 38,909
Videos Min. Date Max. Date Avg. views Avg. likes Avg. comments
COP27 624 2008-02-11 2023-05-12 571,964.35 5,860.32 306.99
IPCC 145 2007-11-03 2023-06-09 350,178.66 8,991.67 1,455.16
C. Refugees 215 2009-11-22 2023-11-06 243,720.27 3,604.81 527.40
Doñana 191 2009-01-11 2023-06-01 101,098.32 3,622.68 246.25
@@
Unlike the Twitter dataset, the period of the comments is longer due to users frequently referencing older videos. Each dataset exhibits peaks at different times based on its unique characteristics. In particular, we observe a strong alignment in the IPCC and COP27 datasets with the event dates (Table 2). Additional insights, including the data collection procedure can be found in Section S1 of the Supplementary Material.
II Results
II.1 Network analysis and structural polarisation
@@
We evaluate the structural polarisation on Twitter by considering the unweighted network of retweets with undirected links indicating a retweet interaction between two users. Previous studies have used retweets as a proxy for the influence between individuals allowing for the computation of their latent ideology [63, 45]. Similarly, we aim to measure how the influence between individuals shapes network organisation. We applied the Metis partition algorithm [68] to split the graph into two clusters, and we calculated the modularity [69], the E-I index [70] and the adaptive E-I index [71] that adjusts for the cluster sizes. These metrics evaluate the topological interactions between users, the mesoscale structure and the extent to which users in different groups interact with one another without considering the content. The results for the four datasets analysed are presented in Fig. 1. Following [34], we computed polarisation metrics in the observed network $\Phi(G)$, as well as in the ensemble of configuration models with equivalent degree distribution $\Phi(G_{CM})$ [72], and calculated the denoised value $\hat{\Phi}(G)=\Phi(G)-\Phi(G_{CM})$. For completeness, we report the standardised values [34] in Fig. S3.
@@
Figure 1: Polarization in the Twitter dataset. Quantification of the polarization in Twitter: (a) modularity, (b) E-I index, and (c) adaptive E-I index. We provide the measures for the observed networks in red, across 100 realisations of the configuration model with preserved degree sequences in blue, and the difference between them (denoised value) in yellow.
@@
All networks exhibit structural polarisation according to the denoised value, suggesting the presence of two groups of strongly connected individuals. In particular, the COP27 and Doñana datasets displayed larger denoised values, suggesting a stronger polarisation around the topics. The higher structural polarisation observed in the COP27 dataset may be attributed to transnational spheres using different languages that do not interact with each other [73]. The first and second largest communities, in terms of the number of messages, are dominated by English. The third and fourth communities primarily consist of Portuguese messages, followed by others where French and Spanish are more prevalent. Further details on this distribution per dataset and community can be found in Section S2.2. The Doñana dataset captures the discussion around a new law related to natural park preservation, which involves political parties and users with strong political affiliations like politicians, journalists, or activists. The lower polarisation observed in the IPCC dataset may be due to its nature as a scientific report release rather than a mainstream international conference like COP, potentially attracting a more scientifically educated audience and inducing a lower level of polarisation. Although the largest community (labelled as 0) is composed mainly of scientists, primarily from the UK, other communities consist of activists (community 3) or politicians (community 6), such as @GretaThunberg and @BarackObama, respectively.
@@
Figure 2: Community interaction network. Interaction between the ten largest communities in (a, b) COP27, (c, d) IPCC, (e, f) Climate Refugees, (g, h) Doñana. We report the interactions via (a, c, e, g) retweets and (b, d, f, h) mentions. Arrows indicate the direction of the interaction, the colour intensity and width, the volume. The size of the dots represents community size. In blue, we show the interaction through retweets and, in orange, through mentions.
While the retweet network has been previously interpreted as the influence between social media users, other types of interactions, such as mentions, often feature different structural properties [61]. We examine the network of retweets and mentions among the $10$ largest communities of the networks in Fig. 2. We identified these communities using modularity optimisation on the directed weighted graph of retweets only [69], and then we incorporate mentions. This approach allows us to identify clusters based on influence and later analyse how they interact through mentions and retweets. The interaction patterns among the largest community (labelled as 0) and the remaining communities vary across datasets and types of interactions. In the COP27 and IPCC networks, the smaller communities mostly retweet content from the largest community, which is more isolated and engages in fewer interactions with the rest of the network. The largest communities include official accounts and major figures, such as the Secretary-General of the United Nations (@antonioguterres) or the Intergovernmental Panel on Climate Change (@IPCC_CH), representing COP27 and the IPCC, respectively. Conversely, the Doñana network exhibits higher flow heterogeneity between communities, with community 5 playing a central role in attracting most of the flow. The detailed overview of the users in the community reveals the presence of the president of Spain (@sanchezcastejon), the ruling political party (@PSOE), and sympathiser journalists (@eldiarioes, @iescolar). There are significant structural differences depending on the type of interaction. For example, in the IPCC network, there are relatively few retweets from community 1, with accounts concentrated flow such as @uksciencechiefUK (Government Office for Science) and @BernieSpofforth (the anti-climate movement figure) to community 0, formed by well-know climate science accounts such as @ed_hawkins or @MrMatthewTodd. However, there is a high volume of mentions between these communities (see details in Section S2.1).
@@
Figure 3: Reliability as a function of political bias in the studied networks. Density plot showing the distribution of reliability as a function of bias for the communities detected in (a) COP27, (b) IPCC, (c) Climate Refugees, and (d) Doñana. For each dataset, the two communities with the largest and most divergent political views are identified. In dark brown the left-biased community and in light brown the right-biased community
II.1.1 Isolation and entropy